OpenStudy (anonymous):

[3f(x)+f(x)g(x)]=36 find f(2) suppose f and g are continuous functions such that g(2)=6

OpenStudy (anonymous):

I don't really know what the continuous function is but i will try to help . in LHS I will take f(x) common so the expression can be written as f(x)[3+g(x)} now g(x)=6 this implies that f(x)[3+6]=36 { from the equation and above } this implies that f(x) = 36/9 this implies f(x) = 4 this implies f(2) = 4 I will be grateful if you can explain me what the continuous function is ?